# Solving Modular Linear Systems with a Constraint by parallel decomposition of the Smith form and extended Euclidean division modulo powers of primes divisors

### v3

@@ 3.4 Using the CRT to solve the modular system
3:     Output Integeral Solutions $x,y$ of the system $Ax-ny=b$.
4:     Form the augmented matrix $[A,-nI]$.
@@
Using the simplification in computation of the invariant factors of $[A,-p^{r}I]$ and decomposition of the problem of solving the modular system over the residues as described in Theorem 4, existence and construction of solutions to Problem 1 can be solved efficiently by parallel computation by using the reduction over moduli $p^{r(p)}$ which are coprime divisors of $n$. Further refinement of the computation is obtained by using efficient arithmetic modulo $q=p^{d}$ where $d$ is the byte length of representation of numbers in digits modulo $p$. Such a computation is described in Appendix A.
@@ 4 Linear system with constraint over a field
Let $\ker A\neq 0$. Next requirement of the problem is that there exists a solution $x$ such that $<w,x>\neq 0$. Hence it is necessary and sufficient along with condition 1 that the system of equations
@@
has a solution set $S_{1}$ which is a strict subset of $S_{0}$. Let $Ax_{0}=b$ then $x_{0}$ belongs to $S_{0}$. If $<w,x_{0}>\neq 0$ then $x_{0}$ does not belong to the solution set $S_{1}$ and this shows that $S_{1}$ is already a strict subset of $S_{0}$. Hence we assume that $<w,x_{0}>=0$. Then the solution set of the above equations is
@@
dim $\dim\mbox{ rowspan }A<\dim\mbox{ rowspan }\left[\begin{array}[]{l}A\\ w^{T}\end{array}\right]$
@@ 4.1 Condition when the linear equation (14) has a unique solution
dim $\dim\mbox{ rowspan }\left[\begin{array}[]{l}A\\ w^{T}\end{array}\right]<\dim\mbox{ rowspan }\left[\begin{array}[]{ll}A&b\\ w^{T}&0\end{array}\right]$
@@ A.1 Solution modulo $q$
Let $s=r-(m+1)$, given the single number
@@
$\begin{array}[]{lclcl}t_{0}&:=&(a_{0}x_{0})\\ t_{0}&=&1\mbox{ \rm{mod} }q\\ \mbox{for $k=1,2,\dots,(t-1)$}&&\\ t_{k}&:=&[t_{(k-1)}]+(a_{0}x_{k}+a_{1}x_{(k-1)}+\ldots+a_{k}x_{0})\\ t_{k}&=&0\mbox{ \rm{mod} }q\\ \mbox{for $j=0,1,\ldots,(t-1)$}&&\\ t_{(t+j)}&:=&[t_{(t-1+j)}]+a_{1}x_{(t-1+j)}+a_{2}x_{(t-2+j)}+\ldots+a_{(t-1)}x_{(1+j)}\\ t_{(t+j)}+y_{j}&=&0\mbox{ \rm{mod} }q\end{array}$ (19)
@@
Proof follows by substitution of $q$-adic representations of $a$, $x$, $y$ in (10) and comparing co-efficients of powers of $q$. While each co-efficient is computed modulo $q$, the carry part which is quotient by $q$ is added to the higher power of $q$ shown as $[t_{(k-1)}]$ in the next value of $t_{k}$ in the iteration. Note that each equation $t_{k}=0$ for $k>0$ has the unknown $x_{k}$, is of the form $a_{0}x_{k}=b_{k}\mbox{ \rm{mod} }q$ where $b_{k}$ depends on previously computed $x_{0},\ldots,x_{(k-1)}$ and only inversion modulo $q$ of $a_{0}$ is required for solving all $x_{k}$. ∎
@@
The iterations (19) require arithmetic modulo $q$ which are byte operations. In a predefined arithmetic library the byte operations are implemented efficiently, for instance by look up tables, for addition and multiplication modulo $q$, inversion modulo $q$. One extra operation required in solving the iterations (19) is the quotient operation by $q$ denoted as $[t_{k}]$ (the carry obtained in evaluating $t_{k}$). Hence modulo $q^{r}$ computations can be carried out efficiently by byte operations which are mod $q$ to solve the Bezout identity and subsequently the Smith form by extending the unimodular matrix computations as shown in Section 4.2.2.

### v4

@@ 3.4 Using the CRT to solve the modular system
3:  Output Integeral Solutions $x,y$ of the system Axny=bAx-ny=b.
4:  Form the augmented matrix [A,nI][A,-nI].
@@
Using the simplification in computation of the invariant factors of $A,-p^{r}I]$ and decomposition of the problem of solving the modular system over the residues as described in Theorem 4, existence and construction of solutions to Problem 1 can be solved efficiently by parallel computation by using the reduction over moduli $p^{r(p)}$ which are coprime divisors of $n$. Further refinement of the computation is obtained by using efficient arithmetic modulo $q=p^{d}$ where $d$ is the byte length of representation of numbers in digits modulo $p$. Such a computation is described in Appendix A.
@@ 4 Linear system with constraint over a field
Let $\ker A\neq 0$. Next requirement of the problem is that there exists a solution $x$ such that $w,x>\neq 0$. Hence it is necessary and sufficient along with condition 1 that the system of equations
@@
has a solution set $S_{1}$ which is a strict subset of $S_{0}$. Let $Ax_{0}=b$ then $x_{0}$ belongs to $S_{0}$. If $w,x_{0}>\neq 0$ then $x_{0}$ does not belong to the solution set $S_{1}$ and this shows that $S_{1}$ is already a strict subset of $S_{0}$. Hence we assume that $w,x_{0}>=0$. Then the solution set of the above equations is
@@
$\dim\mbox{ rowspan }A<\dim\mbox{ rowspan }\left[\begin{array}[]{l}A\\ w^{T}\end{array}\right]$
@@ 4.1 Condition when the linear equation (14) has a unique solution
$\dim\mbox{ rowspan }\left[\begin{array}[]{l}A\\ w^{T}\end{array}\right]<\dim\mbox{ rowspan }\left[\begin{array}[]{ll}A&b\\ w^{T}&0\end{array}\right]$
@@ A.1 Solution modulo $q$
Let s=r(m+1)s=r-(m+1), given the single number
@@
[…]
@@
Proof follows by substitution of $q$-adic representations of $a$, $x$, $y$ in (10) and comparing co-efficients of powers of $q$. While each co-efficient is computed modulo $q$, the carry part which is quotient by $q$ is added to the higher power of $q$ shown as $t_{(k-1)}]$ in the next value of $t_{k}$ in the iteration. Note that each equation $t_{k}=0$ for $k>0$ has the unknown $x_{k}$, is of the form $a_{0}x_{k}=b_{k}\mbox{ \rm{mod} }q$ where $b_{k}$ depends on previously computed $x_{0},\ldots,x_{(k-1)}$ and only inversion modulo $q$ of $a_{0}$ is required for solving all $x_{k}$. ∎
@@
The iterations (19) require arithmetic modulo $q$ which are byte operations. In a predefined arithmetic library the byte operations are implemented efficiently, for instance by look up tables, for addition and multiplication modulo $q$, inversion modulo $q$. One extra operation required in solving the iterations (19) is the quotient operation by $q$ denoted as $t_{k}]$ (the carry obtained in evaluating $t_{k}$). Hence modulo $q^{r}$ computations can be carried out efficiently by byte operations which are mod $q$ to solve the Bezout identity and subsequently the Smith form by extending the unimodular matrix computations as shown in Section 4.2.2.